# Adaptive Teaming in Multi-Drone Pursuit: Simulation, Training, and Deployment

### v1

@@ Abstract
Adaptive teaming, the ability to collaborate with unseen teammates without prior coordination, remains an underexplored challenge in multi-robot collaboration. This paper focuses on adaptive teaming in multi-drone cooperative pursuit, a critical task with real-world applications such as border surveillance, search-and-rescue, and counter-terrorism. We first define and formalize the Adaptive Teaming in Multi-Drone Pursuit (AT-MDP) problem and introduce AT-MDP framework, a comprehensive framework that integrates simulation, algorithm training and real-world deployment. AT-MDP framework provides a flexible experiment configurator and interface for simulation, a distributed training framework with an extensive algorithm zoo (including two newly proposed baseline methods) and an unseen drone zoo for evaluating adaptive teaming, as well as a real-world deployment system that utilizes edge computing and Crazyflie drones. To the best of our knowledge, AT-MDP framework is the first adaptive framework for continuous-action decision-making in complex real-world drone tasks, enabling multiple drones to coordinate effectively with unseen teammates. Extensive experiments in four multi-drone pursuit environments of increasing difficulty confirm the effectiveness of AT-MDP framework, while real-world deployments further validate its feasibility in physical systems. Videos and code are available at https://sites.google.com/view/at-mdp.
Machine Learning, ICML
@@ 1 Introduction
Multi-drone pursuit is an increasingly critical task with wide-ranging real-world applications, such as disaster response, border surveillance, search-and-rescue operations, and many more (Chung et al., 2011; Zhang et al., 2023; Queralta et al., 2020). All of these scenarios are highly dependent on the ability of autonomous drones to coordinate effectively in dynamic environments to pursue and track uncooperative targets. Most existing approaches rely on pre-coordinated strategies, where agents either follow predefined coordination mechanisms, such as conventions, role assignments, and communication protocols, or learn them through extensive interaction over time (Rahman et al., 2024). These methods generally fall into two categories: traditional optimization-based approaches (Shah & Schwager, 2019; Janosov et al., 2017b; Zhou et al., 2016),and reinforcement learning-based approaches (Zhang et al., 2023; Chen et al., 2024; Qi et al., 2024; de Souza et al., 2021; Li et al., 2019; Matignon et al., 2007).
Figure 1: Pre-coordinated teaming: Agents coordinate using predefined mechanisms—such as conventions, role assignment protocols, and communication strategies—or by learning these mechanisms through extensive interaction over time (Rahman et al., 2024). Adaptive teaming: The assumed convention or mechanism is disrupted or fails, requiring adaptation to unseen teammates without prior coordination methods.
However, real-world scenarios are inherently uncertain, and unpredictable, often necessitating rapid adaptation to evolving team compositions. Instead of relying on pre-coordinated teaming, drones must seamlessly integrate with unforeseen teammates in real time. For example, in disaster response, damaged drones may need rapid replacement to maintain search coverage, while in border surveillance, specialized drones might temporarily join to intercept high-priority targets. Adaptive teaming is capable of addressing this critical gap by enabling drones to dynamically adjust their coordination strategies with unforeseen teammates, without relying on prior coordination, as shown in Fig. 1.
To address the challenge of working with unseen drone teammates, we first define and formalize the Adaptive Teaming in Multi-Drone Pursuit (AT-MDP) problem and propose the AT-MDP framework, a comprehensive framework that seamlessly integrates simulation, training, and real-world deployment. Specifically, the simulation component enables flexible customization of multi-drone pursuit scenarios through an environment configurator and provides a realistic simulated environment for training and evaluation via the environment interface. The training component leverages a distributed framework for algorithm learning, incorporating a comprehensive algorithm zoo that includes two newly proposed baseline methods designed for adaptive teaming, along with an unseen drone zoo to assess generalization to diverse teammate behaviors. Finally, the deployment component bridges simulation to real-world applications by integrating a motion capture system and edge nodes, enabling real-time data exchange and decision-making for physical drone coordination.
AT-MDP framework also advances adaptive teaming in machine learning by enabling multiple drones to dynamically collaborate with unseen teammates in complex, continuous-action environments, pushing the boundaries of multi-agent coordination in real-world scenarios. In machine learning, zero-shot coordination (ZSC) (Hu et al., 2020) and ad-hoc teamwork (AHT) (Stone et al., 2010) offer alternative paradigms for handling unseen partners. However, most of the research in this area is confined to training a single learner in simulated 2D video games with discrete action spaces, such as Overcooked (Li et al., 2023; Wang et al., 2024d), Hanabi (Hu et al., 2020; Lucas & Allen, 2022; Canaan et al., 2022; Bard et al., 2020), and Predator-Prey (Barrett et al., 2011; Papoudakis et al., 2021). A recent advancement, NAHT (Wang et al., 2024a), extends AHT to control multiple learners interacting with multiple unseen partners. However, it remains restricted to the SMAC benchmark based on video games (Samvelyan et al., 2019) with discrete actions, limiting its applicability to real-world continuous-action drone tasks.
To validate AT-MDP framework, we conducted extensive experiments in four multi-drone pursuit environments of increasing difficulty, evaluating their adaptability to diverse unseen teammate dynamics. The experiments incorporate multiple unseen drone protocols to ensure a broad spectrum of coordination challenges. The results show that our proposed baseline methods effectively address AT-MDP problem, consistently outperforming existing approaches in coordination success and adaptability. Furthermore, real-world deployment on Crazyflie drones demonstrates the feasibility of AT-MDP framework in physical systems, while highlighting opportunities for further improvements in algorithms, more realistic task scenarios, and enhanced sim-to-real transfer. A case study in a high-complexity scenario further highlights the emergence of adaptive strategies in dynamic, real-world-inspired conditions.
The main contributions of this paper are threefold: (1) To the best of our knowledge, this work is the first to formally define and formulate the adaptive teaming problem in multi-drone pursuit, advancing multi-agent coordination in complex, continuous-action real-world scenarios. (2) We propose a comprehensive AT-MDP framework that integrates flexible simulation, algorithm training, and real-world deployment, paving the way for further research in this area; (3) We validate the effectiveness and feasibility of our baseline methods through extensive simulations and real-world experiments with diverse, previously unseen drone partners.
@@
Table 1: Comparison of AT-MDP framework with main related works. Grey rows represent literature on multi-drone pursuit, while pink rows highlight adaptive teaming studies from the machine learning field. “AT w/o TM” and “AT w/ TM’ denotes adaptive teaming without and with, respectively. A complete tabular comparison of related works is provided in Table 4 in the Appendix.
@@
# Learner # Unseen # Evader Action Space Main Related Task Real-world? AT w/o TM? AT w/ TM?
Voronoi Partitions (Zhou et al., 2016) Multi 0 1 Continuous Pursuit–evasion Game No No No
Bio-pursuit (Janosov et al., 2017b) Multi 0 Multi Continuous Prey–predator Game No No No
M3DDPG (Li et al., 2019) Multi 0 1 Continuous Prey–predator Game No No No
DualCL (Chen et al., 2024) Multi 0 1 Continuous Multi-drone Pursuit No No No
Pursuit-TD3 (de Souza et al., 2021) Multi 0 1 Continuous Multi-drone Pursuit Yes No No
DACOOP-A (Zhang et al., 2023) Multi 0 1 Discrete Multi-drone Pursuit Yes No No
HOLA-Drone (Li et al., 2024a) 1 Multi Multi Continuous Multi-drone Pursuit Yes Yes No
Other-play (Hu et al., 2020) 1 1 0 Discrete Lever Game; Hanabi No Yes No
COLE (Li et al., 2024b) 1 1 0 Discrete Overcooked No Yes No
PLASTIC (Barrett et al., 2017) 1 Multi Multi Discrete Prey-predator Game No No Yes
AATeam (Chen et al., 2020) 1 1 2 Discrete Half Field Offense No No Yes
CIAO (Wang et al., 2024b) 1 Multi Multi Discrete LBF; Wolfpack No No Yes
NAHT (Wang et al., 2024a) Multi Multi Multi Discrete StarCraft2 No No Yes
Our ATMDP Multi Multi Multi Continuous Multi-drone Pursuit Yes Yes Yes
2 Problem Formulation and Related Work
In this section, we first formalize the concept of adaptive teaming in multi-drone pursuit. Next, we discuss the relationship between our work and prior research, as summarized in Table 1. A more detailed literature review and complete tabular comparison of related works can be found in the Appendix A.
@@ 2.1 Problem Formulation
Let ${\mathcal{C}}$ represent the cooperative team, comprising $N$ learners and $M$ uncontrolled teammates. The set of uncontrolled teammates is denoted by ${\mathcal{U}}$. In the multi-drone pursuit task, there exists a set of opponents, denoted as ${\mathcal{E}}$.
@@
Adaptive teaming can be effectively modeled as an extended Adaptive Teaming Decentralized Partially Observable Markov Decision Process (AT-Dec-POMDP). AT-Dec-POMDP is defined by the tuple $({\mathcal{S}},{\mathcal{C}},{\mathcal{A}},{\mathcal{P}},{\mathcal{P}}_{\text{u}},r,{\mathcal{O}},\gamma,T)$, where: where ${\mathcal{S}}$ is the joint state space; ${\mathcal{C}}$ is the set of cooperative agents, including learners (${\mathcal{N}}$) and uncontrolled teammates (${\mathcal{M}}$); ${\mathcal{A}}=\times_{j=1}^{C}{\mathcal{A}}^{j}$ is the joint action space, where $C=N+M$ is the team size; ${\mathcal{P}}_{\text{u}}({\mathcal{M}}|{\mathcal{U}})$ is the uncontrolled teammate sampling function, which defines the probability of sampling a subset ${\mathcal{M}}\subseteq{\mathcal{U}}$ of size $M$ from the set of all uncontrolled teammates ${\mathcal{U}}$; ${\mathcal{P}}(s’|s,a)$ is the transition probability function, representing the probability of transitioning to state $s’\in{\mathcal{S}}$ given the current state $s\in{\mathcal{S}}$ and joint action $a\in{\mathcal{A}}$; $r(s,a)$ is the reward function, representing the team’s reward in state $s$ after taking action $a$; ${\mathcal{O}}$ is the joint observation space, with ${\mathcal{O}}(o|s)$ describing the probability of generating observation $o$ given state $s$; $\gamma\in[0,1]$ is the discount factor; and $T$ is the task horizon.
At each time step $t$, the AT-Dec-POMDP is in state $s_{t}\in{\mathcal{S}}$ and generates a joint observation $o_{t}=(o_{t}^{1},\dots,o_{t}^{C})\sim{\mathcal{O}}(\cdot|s_{t})$. Each agent $j\in{\mathcal{C}}$ observes $o_{t}^{j}$ and maintains its own action-observation trajectory $\tau_{t}^{j}=(o_{0}^{j},a_{0}^{j},\dots,o_{t-1}^{j},a_{t-1}^{j},o_{t}^{j})$. Furthermore, we denote ${\mathcal{T}}^{j}$ as the set of all possible partial observations and action histories for the agent $j$. Using its policy $\pi^{j}$, the agent selects an action $a_{t}^{j}\in{\mathcal{A}}^{j}$. The policies of the $N$ learners, denoted by $\pi^{i}$ for $i\in{\mathcal{N}}$, are learnable.
@@
We define the policy in two approaches: with and without teammate modeling. Adaptive teaming without teammate modeling is closely related to the zero-shot coordination problem (Hu et al., 2020; Carroll et al., 2019), where learners must rapidly coordinate with unseen teammates. Specifically, the policy for a learner $i$ is represented as $\pi^{i}(a_{t}^{i}\mid\tau_{t}^{i})$, where $\tau_{t}^{i}$ denotes the learner’s observation history up to time $t$. On the other hand, adaptive teaming with teammate modeling aligns closely with the ad-hoc teamwork paradigm (Stone et al., 2010), where learners explicitly model their teammates using a separate teammate model $f$. In this case, the policy is defined as $\pi^{i}(a_{t}^{i}\mid\tau_{t}^{i},f(\tau_{t}^{i}))$. The joint action $a_{t}=(a_{t}^{1},\dots,a_{t}^{C})$ determines the next state $s_{t+1}\sim{\mathcal{P}}(s_{t+1}\mid s_{t},a_{t})$, and all agents receive a shared reward $r(s_{t},a_{t})$.
The goal of adaptive teaming is to learn policies $\{\pi^{i}\}_{i\in{\mathcal{N}}}$ that maximize the expected discounted return: ${\mathcal{J}}=\mathbb{E}[R(\tau)]=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\right],$ where $\tau$ denotes the joint trajectory.
Figure 2: Overview of AT-MDP framework. AT-MDP framework consists of three core components: Simulation, Training, and Deployment. The Simulation module provides an Environment Configurator and Interface for flexible scenario customization and a simulated training environment. The Training module operates in a distributed framework, featuring an Algorithm Zoo and an Unseen Drone Zoo for evaluating adaptive teaming. The Deployment module integrates learned policies into real-world tasks using a motion capture system and edge nodes for real-time decision-making, seamlessly bridging simulation, training, and deployment.
2.2 The Relation to Prior Works
In this section, we highlight the key distinctions between AT-MDP framework and prior works. Tables 1 and Table 4 provide a comprehensive comparison across various dimensions, including problem formulation, scope of tasks, and methodological approach. AT-MDP framework introduces a novel adaptive teaming framework tailored for multi-drone pursuit, setting it apart from existing methods. By pushing the boundaries of existing research, AT-MDP framework paves the way for breakthroughs in multi-drone applications for disaster response, border surveillance, and counter-terrorism.
Comparison with Multi-Drone Pursuit Literature (Gray Rows). The studies highlighted in gray focus on pursuit-evasion tasks, with several works (Zhou et al., 2016; Janosov et al., 2017b; Li et al., 2019; Chen et al., 2024) restricted to simulated environments without real-world deployment. Others, such as Pursuit-TD3 (de Souza et al., 2021) and DACOOP-A (Zhang et al., 2023), operate under pre-coordinated settings, where all pursuers are jointly trained to optimize performance in real-world drone pursuit tasks. Unlike these approaches, while HOLA-Drone (Li et al., 2024a) is a recent attempt at adaptive teaming, it is limited to a single learner coordinating with multiple teammates without explicit teammate modeling. AT-MDP framework overcomes these limitations by enabling multiple learners to dynamically adapt and coordinate with unseen teammates in continuous-action real-world multi-drone pursuit tasks.
Comparison with Adaptive Teaming Literature (Pink Rows). In the machine learning community, zero-shot coordination (ZSC) and ad-hoc teamwork (AHT) are closely related to adaptive teaming but have been explored primarily in discrete-action video game environments, as shown in the pink rows of Table 1. Almost all ZSC methods, including Other-play (Carroll et al., 2020), TrajDi (Lupu et al., 2021), MEP (Zhao et al., 2023), LIPO (Charakorn et al., 2023), COLE (Li et al., 2024b), and ZSC-Eval (Wang et al., 2024d), adopt a two-player game formulation with discrete action spaces. For AHT, most existing methods can be interpreted as a specific case of the AT-Dec-POMDP formulation with $N=1$, where a single learner adapts to unseen teammates (Stone et al., 2010; Rahman et al., 2022, 2024; Barrett et al., 2017; Chen et al., 2020; Wang et al., 2024c). The recent NAHT (Wang et al., 2024a) extends AHT to a multi-learner setting, closely resembling the AT-Dec-POMDP formulation with teammate modeling. However, NAHT remains limited to discrete-action decision-making in video game environments, whereas AT-MDP framework advances adaptive teaming by enabling continuous-action decision-making in real-world multi-drone pursuit tasks.
3 The Framework: AT-MDP
To address the adaptive teaming problem in the multi-drone pursuit task, we propose a comprehensive framework called AT-MDP framework, encompassing three key components: Simulation, Training, and Deployment, as illustrated in Fig. 2. The Simulation component allows flexible customization of multi-drone pursuit environments through the environment configurator, while the environment interface provides a simulated environment for training and evaluation. The Training component leverages a distributed training framework to train algorithms through interaction with the environment interface within the simulation. It also includes an algorithm zoo that features a series of baseline methods for solving AT-MDP problem, as well as an unseen drone zoo that serves as a repository of various unseen drone policies to test adaptive teaming capabilities. Finally, the Deployment component facilitates real-world applications using a motion capture system and edge nodes, enabling real-time data exchange and decision-making. Section 3.2 introduces the simulation and deployment modules in detail, while Section 3.1 provides a detailed explanation of the training component.
3.1 Training
The training module comprises three key components: a distributed training framework, an algorithm zoo, and an unseen drone zoo. To enable efficient interaction with environments, AT-MDP framework supports distributed environment sampling, ensuring scalability and adaptability in training processes. The algorithm zoo and the unseen drone zoo are the core components of the training module, providing a series of baseline methods for AT-MDP problem and a diverse population of drone policies for evaluation.
Algorithm Pool. The algorithm pool in AT-MDP framework includes MARL-based pre-coordinated approaches, self-play frameworks, population-based training (PBT) strategies, and our proposed methods for adaptive teaming. For MARL-based approaches, we use the MAPPO algorithm (Yu et al., 2022), a strong baseline designed for multi-drone pursuit tasks, which leverages centralized training with decentralized execution for effective coordination. Self-play and PBT (Carroll et al., 2019) serve as widely used methods for zero-shot coordination, with self-play enabling agents to iteratively learn through interactions with copies of themselves and PBT allowing a population of models to explore various strategies and share knowledge for improved generalization. Both self-play and PBT are implemented using the Independent PPO (Schulman et al., 2017) framework, ensuring scalable and robust training for adaptive teaming tasks.
In addition to existing baseline methods, we propose a novel Open-Ended Population Training (OPT) algorithm to better address AT-MDP problem without teammate modeling. It begins by defining a population of drone strategies, $\Pi=\{\pi_{1},\pi_{2},\cdots,\pi_{n}\}$, where interactions among $C$ teammates are represented as a hypergraph $\mathcal{G}=(\Pi,\mathcal{E},\mathbf{w})$. Here, $\Pi$ represents the set of strategies, $\mathcal{E}$ denotes the interaction relationships, and $\mathbf{w}$ captures the results of these interactions. To evaluate coordination ability, the preference hypergraph $\mathcal{P}\mathcal{G}$ is derived from $\mathcal{G}$, retaining only hyperedges that represent optimal teammate configurations. The cooperative ability of each strategy is quantified by its preference centrality: $\eta_{\Pi}(i)=\frac{d_{\mathcal{P}\mathcal{G}}(i)}{d_{\mathcal{G}}(i)},$ where $d_{\mathcal{P}\mathcal{G}}(i)$ and $d_{\mathcal{G}}(i)$ are the incoming degrees in $\mathcal{P}\mathcal{G}$ and $\mathcal{G}$, respectively. To enhance adaptability, the max-min preference oracle alternates between two steps: identifying the most challenging teammate subsets (minimization) and updating the learner set to maximize performance against these subsets (maximization). The oracle updates the learner set by solving: $\mathcal{N}^{\prime}=\arg\max_{\mathcal{N}}\min_{\mathcal{M}\subseteq\Pi_{-\mathcal{N}}}\phi_{\mathcal{M}}(\mathcal{N}),$ where $\mathcal{N}$ is the learners, $\Pi_{-{\mathcal{N}}}$ the objective $\phi_{\mathcal{M}}(\cdot)$ is formulated using the extended curry operator (Balduzzi et al., 2019), which was originally developed for two-player games. The function $\phi$ evaluates the interactions between $N$ learners and $M$ teammates, producing a real-valued performance outcome. The max-min preference oracle forms the foundation of the OPT algorithm, which evolves the strategy population $\Pi$ and iteratively refines the learner set $\mathcal{N}$ over generations. By dynamically recalibrating the training objectives, OPT promotes robust coordination with diverse, unseen teammates, ensuring adaptability in complex real-world scenarios. For detailed algorithmic steps and implementation, please refer to the Appendix B.
We propose an adaptive teaming with modeling (ATM) baseline to efficiently model unseen drone teammates. ATM incorporates an additional teammate modeling network built on MAPPO (Yu et al., 2022). The modeling network, denoted as $f$, generates team encoding vectors that characterize the strategies of teammates. Specifically, the teammate modeling network $f_{\text{tm}}:{\mathcal{T}}^{i}\rightarrow\mathbb{R}^{n}$, parameterized by $\theta_{tm}$, processes the modeling agent’s history of local observations and actions. The resulting team encoding vector is integrated into the actor network within the MAPPO architecture. For detailed algorithmic implementation, please refer to Appendix F.
Unseen Drone Zoo. To efficiently evaluate adaptive teaming performance, we construct an unseen drone zoo consisting of rule-based methods, traditional optimization-based methods, and learning-based approaches. This pool introduces diverse and challenging teammate behaviors to ensure a robust evaluation of AT-MDP framework in multi-drone pursuit tasks, particularly in continuous-action settings. Specifically, the unseen drone zoo includes three types of algorithms to evaluate adaptive teaming performance. The Greedy Drone follows a rule-based approach, pursuing the closest target while dynamically avoiding obstacles and teammates. The VICSEK Drone, based on the VICSEK algorithm (Janosov et al., 2017a; Zhang et al., 2023; Li et al., 2024a), adopts a bioinspired strategy to mimic swarm-like behaviors, optimizing its path toward the evader while applying repulsive forces to avoid collisions. Finally, the Self-Play Drones use an IPPO-based self-play framework, generating diverse behaviors through randomized training, introducing complexity and variability into the evaluation process. For detailed algorithmic description and implementation, please refer to the Appendix C.

### v2

# AT-Drone: Benchmarking Adaptive Teaming in Multi-Drone Pursuit

@@ Abstract
Adaptive teaming—the capability of agents to effectively collaborate with unfamiliar teammates without prior coordination—is widely explored in virtual video games but overlooked in real-world multi-robot contexts. Yet, such adaptive collaboration is crucial for real-world applications, including border surveillance, search-and-rescue, and counter-terrorism operations. To address this gap, we introduce AT-Drone, the first dedicated benchmark explicitly designed to facilitate comprehensive training and evaluation of adaptive teaming strategies in multi-drone pursuit scenarios. AT-Drone makes the following key contributions: (1) An adaptable simulation environment configurator that enables intuitive and rapid setup of adaptive teaming multi-drone pursuit tasks, including four predefined pursuit environments. (2) A streamlined real-world deployment pipeline that seamlessly translates simulation insights into practical drone evaluations using edge devices and Crazyflie drones. (3) A novel algorithm zoo integrated with a distributed training framework, featuring diverse algorithms explicitly tailored, for the first time, to multi-pursuer and multi-evader settings. (4) Standardized evaluation protocols with newly designed unseen drone zoos, explicitly designed to rigorously assess the performance of adaptive teaming. Comprehensive experimental evaluations across four progressively challenging multi-drone pursuit scenarios confirm AT-Drone’s effectiveness in advancing adaptive teaming research. Real-world drone experiments further validate its practical feasibility and utility for realistic robotic operations. Videos, code and weights are available at https://sites.google.com/view/at-drone.
Keywords: adaptive teaming,multi-robot collaboration,multi-drone pursuit
@@ 1 Introduction
Multi-drone collaboration has become increasingly critical for a variety of real-world applications, including disaster response, border surveillance, and search-and-rescue missions [1, 2, 3]. The success of these missions heavily relies on drones’ capability to effectively coordinate in real-time within dynamic environments with changing team compositions and unpredictable target behaviors. For example, in disaster scenarios, drones may become damaged or depleted during operation, necessitating rapid integration of backup drones to sustain mission effectiveness. Adaptive teaming directly addresses these challenges by enabling drones to dynamically collaborate with previously unseen teammates, significantly enhancing the operational robustness and flexibility of drone fleets.
However, current multi-drone collaboration methodologies predominantly rely on pre-defined coordination mechanisms or extensive prior interactions among drones, thus limiting their adaptability to unexpected or new teammates. Traditional optimization-based approaches [4, 5, 6] and reinforcement learning methods [2, 7, 8, 9, 10, 11] typically utilize fixed conventions, roles, or communication protocols, hindering their performance when encountering unfamiliar teammates or unpredictable environments.
Conversely, existing adaptive teaming research—such as zero-shot coordination (ZSC) [12] and ad-hoc teamwork (AHT) [13]—mainly focuses on simulated environments with discrete action spaces, exemplified by video games such as Overcooked [14, 15], Hanabi [12, 16, 17, 18], and Predator-Prey [19, 20]. Recent advancements like NAHT [21] extend these paradigms to multiple learners but remain restricted within discrete-action domains such as the SMAC benchmark [22], limiting their real-world applicability.
@@
Table 1: Comparison of related work. Grey rows represent literature related to multi-drone pursuit, while pink rows highlight adaptive teaming studies from the machine learning field. “AT w/o TM” and “AT w/ TM” denote adaptive teaming without and with teammate modelling, respectively.
@@
# Learner # Unseen # Opponent Action Space Main Related Task Real-world? AT w/o TM? AT w/ TM?
Voronoi Partitions [6] Multi 0 1 Continuous Pursuit–evasion Game No No No
Bio-pursuit [5] Multi 0 Multi Continuous Prey–predator Game No No No
Uncertainty-pursuit [4] Multi 0 1 Continuous Pursuit–evasion Game No No No
M3DDPG [10] Multi 0 1 Continuous Prey–predator Game No No No
Pursuit-TD3[9] Multi 0 1 Continuous Multi-drone Pursuit Yes No No
DACOOP-A[2] Multi 0 1 Discrete Multi-drone Pursuit Yes No No
GM-TD3 [23] Multi 0 1 Continuous Prey–predator Game No No No
DualCL [7] Multi 0 1 Continuous Multi-drone Pursuit No No No
HOLA-Drone [24] 1 Multi Multi Continuous Multi-drone Pursuit Yes Yes No
Other-play [12] 1 1 0 Discrete Lever Game; Hanabi No Yes No
Overcooked-AI  [25] 1 1 0 Discrete Overcooked No Yes No
TrajDi [26] 1 1 0 Discrete Overcooked No Yes No
MEP [27] 1 1 0 Discrete Overcooked No Yes No
LIPO [28] 1 1 0 Discrete Overcooked No Yes No
HSP [29] 1 1 0 Discrete Overcooked No Yes No
COLE [30] 1 1 0 Discrete Overcooked No Yes No
ZSC-Eval [15] 1 1 0 Discrete Overcooked No Yes No
PLASTIC [31] 1 Multi Multi Discrete Prey-predator Game No No Yes
AATeam [32] 1 1 2 Discrete Half Field Offense No No Yes
LIAM [20] 1 Multi Multi Discrete LBF; Prey-predator Game No No Yes
GPL [33] 1 Multi Multi Discrete LBF; Wolfpack; FortAttack No No Yes
CIAO [34] 1 Multi Multi Discrete LBF; Wolfpack No No Yes
NAHT [21] Multi Multi Multi Discrete StarCraft; MPE No No Yes
As summarized in Table 1, there is a clear lack of benchmarks tailored specifically for studying adaptive teaming in complex, real-world scenarios involving multi-drone collaboration. To address this critical gap, we introduce AT-Drone, the first unified benchmark explicitly designed to integrate adaptive teaming methods from machine learning into practical multi-drone robotics applications. AT-Drone provides a comprehensive and standardized training and evaluation framework, facilitating rapid assessment of adaptive teaming algorithms and effectively bridging theoretical innovations with real-world robotic deployments.
AT-Drone consists of four main components designed to thoroughly study adaptive teaming in multi-drone pursuit tasks: 1. A customizable simulation environment: AT-Drone includes four progressively challenging multi-drone pursuit environments, systematically varying in obstacle complexity, evader numbers, and task difficulty, enabling rigorous testing of adaptive strategies across diverse operational conditions. 2. Streamlined real-world deployment pipeline: AT-Drone integrates practical deployment pipelines that leverage motion capture systems and edge devices (such as Nvidia Jetson Orin Nano). Real-world experiments with Crazyflie drones validate the benchmark’s fidelity and reflect its potential for advancing more complex real-world applications in the future. 3. A novel algorithm zoo: AT-Drone introduces a distributed training infrastructure comprising seven adaptive teaming algorithms adapted and extended from discrete video game environments. To the best of our knowledge, this benchmark represents the first systematic exploration of adaptive teaming strategies tailored to multi-pursuit multi-evader drone pursuit scenarios. 4. Standardized evaluation protocols: AT-Drone provides three distinct “unseen drone zoo” configurations, each demanding unique adaptive collaboration strategies, alongside four specialized evaluation metrics designed to systematically assess algorithm adaptability and robustness.
2 Related Work
As summarised in Table 1, we provide a detailed comparison of related methods across key dimensions, including problem formulation, task scope, and methodological approaches, highlighting the unique positioning of our benchmark within the literature.
Multi-agent pursuit-evasion. Multi-agent pursuit-evasion is closely related to the multi-drone pursuit task. Most existing methods rely on pre-coordinated strategies specifically designed for particular pursuit-evasion scenarios. Traditional approaches often rely on heuristic [5] or optimisation-based strategies [6, 4]. In recent years, deep reinforcement learning (DRL) has been widely adopted for pre-coordinated multi-drone pursuit tasks. M3DDPG [10] and GM-TD3 [23] extend standard DRL algorithms, such as TD3 [35] and DDPG [36], specifically for multi-agent pursuit in simulated environments. Pursuit-TD3 [9] applies the TD3 algorithm to pursue a target with multiple homogeneous agents, which is validated through both simulations and real-world drone demonstrations. Zhang et al. [2] introduces DACOOP-A, a cooperative pursuit algorithm that enhances reinforcement learning with artificial potential fields and attention mechanisms, which is evaluated in real-world drone systems. DualCL [7] addresses multi-UAV pursuit-evasion in diverse environments and demonstrates zero-shot transfer capabilities to unseen scenarios, though only in simulation. The most recent work, HOLA-Drone [24], claims to be the first ZSC framework for multi-drone pursuit. However, it is limited to controlling a single learner, restricting its applicability to broader multi-agent settings.
Adaptive Teaming. The adaptive teaming paradigm can be broadly categorised into two aspects: adaptive teaming without teammate modelling (AT w/o TM) and adaptive teaming with teammate modelling (AT w/ TM), which correspond to the zero-shot coordination (ZSC) and ad-hoc teamwork (AHT) problems in the machine learning community, respectively. AT w/o TM focuses on enabling agents to coordinate with unseen teammates without explicitly modelling their behaviours. Other-Play [12] introduces an approach that leverages symmetries in the environment to train robust coordination policies, applied to discrete-action tasks like the Lever Game and Hanabi. Similarly, methods such as Overcooked-AI [25], TrajDi [26], MEP [27], LIPO [28], and ZSC-Eval [15] study collaborative behaviours in Overcooked, where agents learn generalisable coordination strategies with diverse unseen partners. While these approaches demonstrate promising results, they are limited to single-learner frameworks in simplified, discrete-action domains like Overcooked and Hanabi. They lack scalability to multi-agent settings, continuous action spaces, and the complexities of real-world applications. AT w/ TM, on the other hand, explicitly models the behaviour of unseen teammates to facilitate effective collaboration. Early methods like PLASTIC [31] reuse knowledge from previous teammates or expert input to adapt to new teammates efficiently. AaTeam [32] introduces attention-based neural networks to dynamically process and respond to teammates’ behaviours in real-time. More advanced approaches, such as LIAM [20], employ encoder-decoder architectures to model teammates using local information from the controlled agent. GPL [33] and CIAO [34] leverage GNNs to address the challenges of dynamic team sizes in AHT. Extending from the AHT settings, NAHT [21] enables multiple learners to collaborate and interact with diverse unseen partners in N-agent scenarios. Despite their progress, these methods remain confined to discrete action spaces and simulated benchmarks, limiting their applicability to real-world, continuous-action tasks.
Figure 1: Overview of the AT-Drone Benchmark, comprising four key components: (I) a customizable simulation environment featuring varied multi-drone pursuit tasks with adjustable complexity; (II) a streamlined real-world deployment pipeline employing motion capture systems and edge devices to facilitate realistic drone validation; (III) a distributed training framework equipped with diverse adaptive teaming algorithms for multi-drone pursuit task; and (IV) standardized evaluation protocols, leveraging diverse unseen teammate configurations to rigorously evaluate adaptive teaming performance and robustness across distinct strategies.
3 The Benchmark: AT-Drone
@@ 3.1 Problem Formulation
Let ${\mathcal{C}}$ represent the cooperative team, comprising $N$ learners and $M$ uncontrolled teammates. The set of uncontrolled teammates is denoted by ${\mathcal{U}}$. In the multi-drone pursuit task, there exists a set of opponents, denoted as ${\mathcal{E}}$. Adaptive teaming can be effectively modeled as an extended Adaptive Teaming Decentralized Partially Observable Markov Decision Process (AT-Dec-POMDP). AT-Dec-POMDP is defined by the tuple $({\mathcal{S}},{\mathcal{C}},{\mathcal{A}},{\mathcal{P}},r,{\mathcal{O}},\gamma,T)$, where: where ${\mathcal{S}}$ is the joint state space; ${\mathcal{C}}$ denotes the set of cooperative agents, consisting of learners (${\mathcal{N}}$) and uncontrolled teammates (${\mathcal{M}}$), where ${\mathcal{M}}$ is sampled according to ${\mathcal{P}}_{\text{u}}({\mathcal{M}}|{\mathcal{U}})$ from the complete set of uncontrolled teammates ${\mathcal{U}}$; ${\mathcal{A}}=\times_{j=1}^{C}{\mathcal{A}}^{j}$ is the joint action space, where $C=N+M$ is the team size; ${\mathcal{P}}(s’|s,a)$ is the transition probability function, representing the probability of transitioning to state $s’\in{\mathcal{S}}$ given the current state $s\in{\mathcal{S}}$ and joint action $a\in{\mathcal{A}}$; $r(s,a)$ is the reward function, representing the team’s reward in state $s$ after taking action $a$; ${\mathcal{O}}$ is the joint observation space, with ${\mathcal{O}}(o|s)$ describing the probability of generating observation $o$ given state $s$; $\gamma\in[0,1]$ is the discount factor; and $T$ is the task horizon.
@@
Additionally, we denote the policy of agent $j$ as $\pi^{j}$, through which the agent selects an action $a_{t}^{j}\in{\mathcal{A}}^{j}$, and the policies of the $N$ learners ($\pi^{i}$, for $i\in{\mathcal{N}}$) are learnable; we consider two approaches for defining these policies: with and without teammate modeling. Adaptive teaming without teammate modeling is closely related to the zero-shot coordination problem [12, 25], where learners could coordinate with unseen teammates. Specifically, the policy for a learner $i$ is represented as $\pi^{i}(a_{t}^{i}\mid\tau_{t}^{i})$, where $\tau_{t}^{i}$ denotes the learner’s observation history up to time $t$. On the other hand, adaptive teaming with teammate modeling aligns closely with the ad-hoc teamwork paradigm [13], where agents could coordinate with previously unknown teammates by explicitly modeling their behavior and characteristics. In this case, the policy is defined as $\pi^{i}(a_{t}^{i}\mid\tau_{t}^{i},f(\tau_{t}^{i}))$. The joint action $a_{t}=(a_{t}^{1},\dots,a_{t}^{C})$ determines the next state $s_{t+1}\sim{\mathcal{P}}(s_{t+1}\mid s_{t},a_{t})$, and all agents receive a shared reward $r(s_{t},a_{t})$. The goal of adaptive teaming is to learn policies $\{\pi^{i}\}_{i\in{\mathcal{N}}}$ that maximize the expected discounted return: ${\mathcal{J}}=\mathbb{E}[R(\tau)]=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\right],$ where $\tau$ denotes the joint trajectory.